# AP-RASC-2025.bib

### v1

@@ 1 Introduction
RFI detection for contemporary and future radio telescopes is becoming a major challenge. With greater sensitivity and data volumes, observatories must balance rising data-processing requirements while addressing increasingly pervasive RFI sources and subject to operational constraints. Spiking Neural Networks (SNNs) are a compelling approach towards real-time RFI detection due to their inherently time-varying nature and extreme energy efficiency when deployed on neuromorphic hardware [davies_loihi_2018, bos_sub-mw_2022]. However, unlike Artificial Neural Networks (ANNs), which communicate through weighted continuous and instantaneous values, SNNs communicate through strictly binary but asynchronous ‘spikes’. Encoding and decoding information for use in SNNs is challenging, but if successful allows th=e network to exploit continuous time dynamics to process spatio-temporal information efficiently. Building on recent advances in SNN-based RFI detection [pritchard_rfi_2024, pritchard_supervised_2024, pritchard_spiking_2024], this work investigates including polarisation information and further presents estimates of power-consumption when deployed on neuromorphic hardware. Using a simulated dataset derived from the Hydrogen Epoch of Reionisation Array (HERA) simulator, we demonstrate state-of-the-art detection performance and compelling operational estimates using SynSense’s Xylo hardware as a model deployment platform.
@@ 2 RFI Detection with SNNs
We utilise a known supervised time-series segmentation approach for RFI detection [pritchard_supervised_2024] modified to include multiple polarisation channels:
@@
We consider latency encoding specifically, which has previously been shown to be an effective spike encoding for this task [pritchard_supervised_2024] and two methods to include polarisation information, directly including all four polarisations as separate input channels (XX, XY, YX and YY) and a Degree of Polarisation (DoP) metric [hashimoto_rfi_2024]. Moreover, we investigate the effect of a divisive normalisation technique inspired from neuroscience research that has been shown to be helpful in this task [pritchard_spiking_2024]. Divisive normalisation is a pre-processing technique applied before the calculation of the DoP metric. The HERA dataset has been well-described elsewhere [mesarcik_learning_2022, pritchard_supervised_2024] and we prepare a similar dataset but retain all polarisation channels, which are openly available online [^1] https://doi.org/10.5281/zenodo.14676274. In addition to training SNNs on the full spectrogram size, like prior works [mesarcik_learning_2022, pritchard_rfi_2024, pritchard_spiking_2024, pritchard_supervised_2024] we divide the original spectrograms into smaller patches for more manageable training, and in this work specifically, to conceivably fit on contemporary neuromorphic hardware.
We employed the snnTorch framework for SNN training [eshraghian_training_2023], with data-driven parallelism managed by the Lightning library [falcon_pytorch_2019]. The spiking neuron model is a first-order Leaky integrate and Fire (LiF) neuron. The Xylo platform is intended for edge-inference of SNNs, it supports 16 input channels, 1000 internal neurons and 16 output channels [noauthor_xylo-audio_2022]. We trained SNNs including full-polarisation information, DoP information, both with and without divisive-normalisation and then for full spectrograms, $16\times 16$ spectrogram patches and smaller $15\times 16$ patches where including DoP and $2\times 16$ patches where including all polarisations. The network is otherwise a multi-layer feed-forward network of spiking neurons. Table 1 summarises the model input and output sizes for each configuration tested. In all cases, the number of hidden neurons is 512 and the final layer contains either 16, 4 or 8 neurons in the case of no, full or DoP polarisation information respectively.
@@
We measure detection performance with respect to per-pixel accuracy, Area Under the Receiver Operating Curve (AUROC), Area Under the Precision Recall Curve (AUPRC) and F1-Score, focusing on the class-balanced metrics AUPRC and F1-Score in particular, in line with other literature [mesarcik_learning_2022, pritchard_spiking_2024].
@@ 3 Computational Power Requirement Modelling
A key hindrance to including machine-learning-based methods in operational radio astronomy observation is the computational requirements, which are significantly guided by energy usage. We provide three energy usage estimates: the first is a lower-bound based on the spiking activity of the trained network; the second is an upper-bound based on the maximum power draw of Xylo hardware ($550\mu W$), and a final estimate combining the measured idle power-draw of the Xylo hardware $216\mu W$ [bos_sub-mw_2022] and the lower-bound estimate.
We estimate a proxy-number of FLOPs per layer $l$, based on known literature[barchid_spiking_2023] as:
@@
where $C_{in}$ is the number of channels in, $C_{out}$ is the number channels out and $R_{s}(l)$ is the spiking rate of layer $l$. We measure the average spike rate of the entire network over the test dataset post-training to provide an indicative measurement. We calculate a lower-bound power usage estimate by multiplying $FLOPs_{SNN}$ by $0.9pJ$, the power usage of a single add operation at 40nm CMOS technology[horowitz_11_2014].
@@ 4 Results
[mesarcik_learning_2022] AOFlagger 0.973 0.880 0.873
[mesarcik_learning_2022] Auto-Encoder 0.973 0.880 0.873
[pritchard_rfi_2024] Auto-Encoder 0.994 0.959 0.945
[pritchard_rfi_2024] ANN2SNN 0.944 0.910 0.953
[pritchard_supervised_2024] BPTT 0.929 0.785 0.761
[pritchard_spiking_2024] BPTT + DN 0.968 0.914 0.907
@@ Acknowledgements
[…]

### v2

# Polarisation-Inclusive Spiking Neural Networks for Real-Time RFI Detection in Modern Radio Telescopes

@@ 1 Introduction
RFI detection for contemporary and future radio telescopes is becoming a major challenge. With greater sensitivity and data volumes, observatories must balance rising data-processing requirements while addressing increasingly pervasive RFI sources and subject to operational constraints. Spiking Neural Networks (SNNs) are a compelling approach towards real-time RFI detection due to their inherently time-varying nature and extreme energy efficiency when deployed on neuromorphic hardware [1, 2]. However, unlike Artificial Neural Networks (ANNs), which communicate through weighted continuous and instantaneous values, SNNs communicate through strictly binary but asynchronous ‘spikes’. Encoding and decoding information for use in SNNs is challenging, but if successful allows th=e network to exploit continuous time dynamics to process spatio-temporal information efficiently. Building on recent advances in SNN-based RFI detection [3, 4, 5], this work investigates including polarisation information and further presents estimates of power-consumption when deployed on neuromorphic hardware. Using a simulated dataset derived from the Hydrogen Epoch of Reionisation Array (HERA) simulator, we demonstrate state-of-the-art detection performance and compelling operational estimates using SynSense’s Xylo hardware as a model deployment platform.
@@ 2 RFI Detection with SNNs
We utilise a known supervised time-series segmentation approach for RFI detection [4] modified to include multiple polarisation channels:
@@
We consider latency encoding specifically, which has previously been shown to be an effective spike encoding for this task [4] and two methods to include polarisation information, directly including all four polarisations as separate input channels (XX, XY, YX and YY) and a Degree of Polarisation (DoP) metric [6]. Moreover, we investigate the effect of a divisive normalisation technique inspired from neuroscience research that has been shown to be helpful in this task [5]. Divisive normalisation is a pre-processing technique applied before the calculation of the DoP metric. The HERA dataset has been well-described elsewhere [7, 4] and we prepare a similar dataset but retain all polarisation channels, which are openly available online [^1] https://doi.org/10.5281/zenodo.14676274. In addition to training SNNs on the full spectrogram size, like prior works [7, 3, 5, 4] we divide the original spectrograms into smaller patches for more manageable training, and in this work specifically, to conceivably fit on contemporary neuromorphic hardware.
We employed the snnTorch framework for SNN training [8], with data-driven parallelism managed by the Lightning library [9]. The spiking neuron model is a first-order Leaky integrate and Fire (LiF) neuron. The Xylo platform is intended for edge-inference of SNNs, it supports 16 input channels, 1000 internal neurons and 16 output channels [10]. We trained SNNs including full-polarisation information, DoP information, both with and without divisive-normalisation and then for full spectrograms, $16\times 16$ spectrogram patches and smaller $15\times 16$ patches where including DoP and $2\times 16$ patches where including all polarisations. The network is otherwise a multi-layer feed-forward network of spiking neurons. Table 1 summarises the model input and output sizes for each configuration tested. In all cases, the number of hidden neurons is 512 and the final layer contains either 16, 4 or 8 neurons in the case of no, full or DoP polarisation information respectively.
@@
We measure detection performance with respect to per-pixel accuracy, Area Under the Receiver Operating Curve (AUROC), Area Under the Precision Recall Curve (AUPRC) and F1-Score, focusing on the class-balanced metrics AUPRC and F1-Score in particular, in line with other literature [7, 5].
@@ 3 Computational Power Requirement Modelling
A key hindrance to including machine-learning-based methods in operational radio astronomy observation is the computational requirements, which are significantly guided by energy usage. We provide three energy usage estimates: the first is a lower-bound based on the spiking activity of the trained network; the second is an upper-bound based on the maximum power draw of Xylo hardware ($550\mu W$), and a final estimate combining the measured idle power-draw of the Xylo hardware $216\mu W$ [2] and the lower-bound estimate.
We estimate a proxy-number of FLOPs per layer $l$, based on known literature[11] as:
@@
where $C_{in}$ is the number of channels in, $C_{out}$ is the number channels out and $R_{s}(l)$ is the spiking rate of layer $l$. We measure the average spike rate of the entire network over the test dataset post-training to provide an indicative measurement. We calculate a lower-bound power usage estimate by multiplying $FLOPs_{SNN}$ by $0.9pJ$, the power usage of a single add operation at 40nm CMOS technology[12].
@@ 4 Results
[7] AOFlagger 0.973 0.880 0.873
[7] Auto-Encoder 0.981 0.927 0.910
[3] Auto-Encoder 0.994 0.959 0.945
[3] ANN2SNN 0.944 0.910 0.953
[4] BPTT 0.929 0.785 0.761
[5] BPTT + DN 0.968 0.914 0.907
@@ Acknowledgements
This work was supported by a Westpac Future Leaders Scholarship, an Australian Government Research Training Program Fees Offset and an Australian Government Research Training Program Stipend. This work was supported by resources provided by the Pawsey Supercomputing Centre with funding from the Australian Government and the Government of Western Australia.